The temperature of 200 g of water decreases from 20 °C to 0 °C in 10 minutes. Calculate

i. the heat lost by the water.

ii. the rate at which the water is losing heat.

iii. the time taken for the water at 0 °C to turn completely into ice.


1
Expert's answer
2022-04-11T12:02:23-0400

1.

"Q=cm\u2206T=16.8~kJ,"

2.

"\dot Q=\frac Qt=28~\frac Js,"

3.

"Q_1=cm∆T+\lambda m=83.6~kJ,"

"t_1=\frac{Q_1}{\dot Q}=49.8~\text{min}."